Diffusion Morning Report 02/16/05

Ted Berg tberg at snf.stanford.edu
Wed Feb 16 06:37:13 PST 2005

Good Morning All,
    Looking Good this morning.  Epi HCl cylinder changed yesterday will 
be back on near noon. SiGe-gas valve problems parts ordered.
Thermco poly and TEOS2 testing.  Ebeam is in High vac.  Everything else 
is OK for now.

  	Date: 	2/16/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	Up 	  	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	down 	Qual testing 	 
Tylan 8-SiGe 	down 	gas flow issues 	0/30
Tylan 9-Poly 	Up 	  	20.69/30
Tylan 10-Nitride 	Up 	  	16.95/30
Tylan 11-BPSG-LTO 	Up 	  	23.21/30
Tylan 12-FGA 	Up 	  	 
Tystar 	yellow 	  	6.93/30
Doped Poly 	GONE 	  	 
TEOS 	up 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
SiH4 for furnaces 	  	Auto changeover working well 	A on line 380 PSI

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050216/f102fd79/attachment.html>

More information about the diffusion mailing list